# Reward-Robust RLHF in LLMs

### v1

@@ B.1 Training Pipeline
The training process is divided into two stages. In the first stage, we train a normal one-head reward model following the loss function in (3). In the second stage, we leverage a MSE loss to train the reward model, which is first introduced in Wu et al. (2024). The loss function for a single head $i$ is given by:
$\ell_{i}=\left\{r_{i}^{+}-\alpha\left[\widehat{p}(a^{+}\succ a^{-})-\frac{1}{2}\right]\right\}^{2}+\left\{r_{i}^{-}-\alpha\left[\widehat{p}(a^{+}\succ a^{-})-\frac{1}{2}\right]\right\}^{2},$ (8)
where $\widehat{p}(a^{+}\succ a^{-})$ is derived from a separate Bradley-Terry model, as defined in (2), and is the output of the normal reward model trained in the first stage. The reward is modeled as a Gaussian distribution, with each head $i$ producing two outputs: one representing the mean and the other representing the standard deviation. A sample from this distribution is then output as the reward. We employ the reparameterization handle the non-differentiability of the sampling process:
$r_{i}^{+}=\mu_{i}^{+}+a\cdot\sigma_{i}^{+},\quad r_{i}^{-}=\mu_{i}^{-}+a\cdot\sigma_{i}^{-}$
@@
Why the standard deviation reflects the confidence of the head?
The use of MSE loss to train the ensembled reward model ensures that the output’s standard deviation reflects the confidence of the model. To understand this, we compute the gradient of the MSE loss with respect to the standard deviation outputs $\sigma_{i}^{+}$ and $\sigma_{i}^{-}$, where $k=\widehat{p}(a^{+}\succ a^{-})-\frac{1}{2}$ is a constant,
@@
A positive gradient indicates that with more optimization steps, the output standard deviation $\sigma_{i}^{+}$ (or $\sigma_{i}^{-}$) decreases, reflecting higher confidence from the reward head in its scoring.
@@ B.3 Experimental results of BRME
Another important performance measurement is the reward margin between chosen and rejected responses. A larger reward margin indicates a greater ability of the reward model to differentiate between better and worse responses, which is critical for guiding the optimization process effectively. We provide a reward margin comparison between BRME and the traditional reward model in Figure 8. The results show that the reward margin in BRME is significantly larger than in the traditional reward model, suggesting that BRME has a stronger capability to distinguish between high-quality and low-quality responses. This improved differentiation leads to more accurate guidance during training, helping the model focus on better responses more consistently. The increased reward margin also reduces the likelihood of reward hacking, as the model is less likely to be misled by small differences between responses. Overall, the larger reward margin in BRME demonstrates its advantage in promoting better alignment with human preferences and improving the robustness of the training process.
@@ Appendix C Additional Experiments
Moreover, in tasks like ANLI-r1, which initially favored standard RLHF at 200 steps, the performance of $\lambda=0.4$ surpasses the nominal reward strategy by the 800-step mark. This indicates that while the nominal reward model may provide faster short-term gains, incorporating robustness into the optimization process enables more consistent long-term improvements. Similarly, on the ARC-challenge, $\lambda=0.6$ outperforms standard RLHF with a 1.35% gain, further confirming the utility of the reward-robust approach for challenging tasks where the reward model may struggle with accuracy in the initial stages.
@@
On the other hand, the conservative nature of the reward-robust RLHF framework, mitigates this risk. By accounting for uncertainty in the reward model, the framework effectively narrows the optimization space, allowing the model to avoid over-optimization based on potentially erroneous reward signals. This results in more stable, long-term gains in performance, as evidenced by the consistent improvements across a range of tasks.
In summary, while standard RLHF may achieve faster short-term improvements in some cases, the reward-robust RLHF framework proves to be more reliable over longer training periods. The inclusion of robustness not only enhances performance but also stabilizes the training process, making it more resilient to the inherent imperfections of reward models, especially in complex and challenging tasks.
@@
Table 1: Accuracy improvement for each $\lambda$ on all the benchmarks tested at 200 step in PPO. Baseline here is the SFT model, which is LLaMa3-8B-Instruct Dubey et al. (2024).
@@
Table 2: Accuracy improvement for each $\lambda$ on all the benchmarks tested at 800 step in PPO. Baseline here is the SFT model, which is LLaMa3-8B-Instruct Dubey et al. (2024).
@@
Overall, the comparison shows that while the over-scoring strategy can occasionally provide short-term performance boosts, particularly in context-heavy tasks, the under-scoring (minimum reward) setting generally yields more reliable and stable improvements across a wider range of benchmarks. This underscores the value of conservative reward modeling, especially when dealing with complex reasoning tasks or when the reward model itself may be noisy or imperfect. The results further validate the robustness of the reward-robust RLHF approach, where minimizing the reward signal helps mitigate the risk of performance degradation over time.
@@
C.3 Ablation study on reward model
@@
In this section, we present ablation study results on the reward models to highlight the superiority of BRME over the traditional RM. The coverage and margin advantages are detailed in Appendix B.3. Here, we provide a comparison of accuracy and the direct effect on the PPO process.
We first compared the accuracy of the single-head reward model trained with MLE loss to the nominal head of the Bayesian ensemble reward model (BRME) on the preference dataset. The results in Table 5 clearly demonstrate that BRME achieves higher accuracy across most benchmarks, indicating a better ability to distinguish between preferred and non-preferred responses. Specifically, BRME outperforms the traditional RM by a notable margin across different categories such as General QA, Writing, Comprehension, and Math. The overall accuracy of BRME reaches 75.6%, compared to 74.5% for the traditional RM, showing that BRME’s ability to model uncertainty provides a tangible advantage in distinguishing between correct and incorrect responses.
@@
We also examined the impact of these reward models on the PPO process by comparing the performance of actors trained using BRME and the traditional RM. The results indicate that actors trained with BRME benefit from more stable and reliable reward signals, leading to smoother training curves and improved final performance. This is particularly evident in tasks requiring more nuanced decision-making, where BRME’s broader reward distribution prevents the actor from overfitting to narrow or incorrect reward signals.
@@ C.4 Ablation study on integration strategy
The mean strategy, while easier to implement, lacks the ability to differentiate between high-quality and low-quality reward signals effectively. By averaging all signals, it fails to account for the underlying uncertainty or variance in the reward model. In contrast, the reward-robust RLHF approach, particularly with $\lambda=0.6$, preserves the benefits of more robust exploration and ensures that the model can continue improving in the later stages of training, as observed in the experiments.
@@
Figure 9: Ablation study on reward model and integration strategy.
Appendix D Dataset Descriptions
Table 6 outlines the various datasets used for training BRME and the PPO models. These datasets span a wide range of task types, ensuring that the reward model is exposed to diverse examples during training, helping it generalize across different domains. In UltraFeedBack, we selected response pairs where the score is greater than or equal to 2 to form the chosen and rejected response pairs, which were then used to train the BRME. By focusing on high-quality response pairs, we aim to ensure that the reward model is robust and capable of accurately distinguishing between better and worse responses.
For evaluation, we utilized a comprehensive set of benchmarks, as detailed in Table 7. These benchmarks cover various domains, including general knowledge, logical reasoning, commonsense, and mathematical reasoning, allowing us to thoroughly assess the performance of the reward models across different types of tasks. To ensure consistency and reliability in our evaluation process, we employed the LM-evaluation-harness framework, which is a widely-used standard for evaluating LLMs Gao et al. (2024). This framework provides a standardized and rigorous approach to comparing model performance across a diverse set of tasks, ensuring that the results are comparable with other works in the field.
By using a diverse set of training datasets and evaluation benchmarks, we aim to provide a comprehensive assessment of our model’s capabilities. The combination of varied training data ensures the robustness of the reward model, while the evaluation benchmarks test the model’s ability to generalize across different domains, making our findings relevant to a wide range of applications.
Table 6: Description of the datasets in the training pipeline.
Datasets Size Description
Table 7: Description of the benchmarks.
Benchmark Size Description
TruthfulQA A QA task aimed at evaluating the truthfulness and factual accuracy of model responses.
MMLU Knowledge-based multi-subject multiple choice questions for academic evaluation.
GSM8K A benchmark of grade school math problems aimed at evaluating reasoning capabilities.
FDA Tasks for extracting key-value pairs from FDA documents to test information extraction.
Arithmetic A small battery of 10 tests that involve asking language models a simple arithmetic problem in natural language.
ANLI Adversarial natural language inference tasks designed to test model robustness. It collected via an iterative, adversarial human-and-model-in-the-loop procedure. It consists of three rounds that progressively increase in difficulty and complexity.
@@ Appendix E Proof for Lemma 1
Let the reward model provide a constant reward $r(s,a)=c$ for all actions $a\in\mathcal{A}$ and states $s\in\mathcal{S}$ during PPO training. Then, the gradient of the PPO objective function with respect to the policy parameters $\theta$ is zero, implying that the actor cannot be optimized under such a reward model.
@@
If the reward model provides a constant reward $r(s_{t},a_{t})=c$ for all actions, the state-action value function and the state value function become:
@@ Appendix F Future works: Heterologous reward fusion
Heterologous Reward Fusion (HRF) aims to enhance the robustness and coverage of the uncertainty set in our reward models by incorporating multiple heterologous reward sources. This method involves combining several different reward models trained on diverse datasets, such as Baichuan2-33B Yang et al. (2023), Qwen2-72B Yang et al. (2024), and LLaMa3-8B Dubey et al. (2024). The key advantage of using heterologous models lies in the diversity of their training data, which produces more varied reward scores and helps to enrich the uncertainty set.
@@
However, one of the primary challenges in heterologous reward fusion is that each reward source has different value ranges, making direct comparisons potentially unfair. To address this, we perform empirical reward normalization. We score a separate dataset, HH-RLHF Bai et al. (2022a), using each reward source and compute the mean and variance for each. During PPO training, we transform the rewards from each source by normalizing them based on their respective means and standard deviations. This normalization helps ensure that the reward signals from different models are comparable, allowing for a more fair integration of the rewards.
Table 8 presents the results of our heterologous reward fusion experiment, showing the reward distribution characteristics (max, min, mean, and standard deviation) for each model, along with their accuracy. Accuracy here is defined as the proportion of instances where the model correctly identifies the chosen response as superior to the rejected one.
However, it is important to note that we did not conduct end-to-end PPO experiments incorporating the full reward model pipeline in this exploration. The primary reason for this limitation is the significant computational resources and time required to carry out such experiments comprehensively. Running full-scale PPO experiments, especially when integrating multiple heterologous reward models, is computationally expensive and requires extended periods of training, particularly when dealing with large models like Baichuan2-33B and Qwen2-72B. Moving forward, our next steps include performing the full end-to-end PPO experiments to evaluate the impact of HRF on the performance of the trained policy.
Table 8: Heterologous Reward Model.

### v3

@@ B.1 Training Pipeline
The training process is divided into two stages. In the first stage, we train a normal one-head RM following the loss function in (3). In the second stage, we leverage a MSE loss to train the RM, which is first introduced in (Wu et al., 2024). The loss function for a single head $i$ is given by:
$\ell_{i}=\left\{r_{i}^{+}-\alpha\left[\widehat{p}(a^{+}\succ a^{-})-\frac{1}{2}\right]\right\}^{2}+\left\{r_{i}^{-}-\alpha\left[\widehat{p}(a^{+}\prec a^{-})-\frac{1}{2}\right]\right\}^{2},$ (8)
where $\widehat{p}(a^{+}\succ a^{-})$ is derived from a separate Bradley-Terry model, as defined in Eq. (2), and is the output of the normal RM trained in the first stage. The reward is modeled as a Gaussian distribution, with each head $i$ producing two outputs: one representing the mean and the other representing the std. A sample from this distribution is then output as the reward. We employ the reparameterization handle the non-differentiability of the sampling process:
$r_{i}^{+}=\mu_{i}^{+}+a\cdot\sigma_{i}^{+},\quad r_{i}^{-}=\mu_{i}^{-}+a\cdot\sigma_{i}^{-},$
@@
Why the std reflects the confidence of the head?
The use of MSE loss to train the BRME ensures that the output’s std reflects the confidence of the model. To understand this, we compute the gradient of the MSE loss with respect to the std outputs $\sigma_{i}^{+}$ and $\sigma_{i}^{-}$, where $k=\widehat{p}(a^{+}\succ a^{-})-\frac{1}{2}$ is a constant,
@@
A positive gradient indicates that with more optimization steps, the output std $\sigma_{i}^{+}$ (or $\sigma_{i}^{-}$) decreases, reflecting higher confidence from the reward head in its scoring.
@@ B.3 Experimental results of BRME
Another important performance measurement is the reward margin between chosen and rejected responses. A larger reward margin indicates a greater ability of the RM to differentiate between better and worse responses, which is critical for guiding the optimization process effectively. We provide a reward margin comparison between BRME and the traditional RM in Figure 6. The results show that the reward margin in BRME is significantly larger than in the traditional RM, suggesting that BRME has a stronger capability to distinguish between high-quality and low-quality responses. This improved differentiation leads to more accurate guidance during training, helping the model focus on better responses more consistently. The increased reward margin also reduces the likelihood of reward hacking, as the model is less likely to be misled by small differences between responses. Overall, the larger reward margin in BRME demonstrates its advantage in promoting better alignment with human preferences and improving the robustness of the training process.
@@ Appendix C Additional Experiments
Moreover, in tasks like ANLI-r1, which initially favored standard RLHF at 200 steps, the performance of $\lambda=0.4$ surpasses the nominal reward strategy by the 800-step mark. This indicates that while the nominal RM may provide faster short-term gains, incorporating robustness into the optimization process enables more consistent long-term improvements. Similarly, on the ARC-challenge, $\lambda=0.6$ outperforms standard RLHF with a 1.35% gain, further confirming the utility of the reward-robust approach for challenging tasks where the RM may struggle with accuracy in the initial stages.
@@
On the other hand, the conservative nature of the reward-robust RLHF framework, mitigates this risk. By accounting for uncertainty in the RM, the framework effectively narrows the optimization space, allowing the model to avoid over-optimization based on potentially erroneous reward signals. This results in more stable, long-term gains in performance, as evidenced by the consistent improvements across a range of tasks.
In summary, while standard RLHF may achieve faster short-term improvements in some cases, the reward-robust RLHF framework proves to be more reliable over longer training periods. The inclusion of robustness not only enhances performance but also stabilizes the training process, making it more resilient to the inherent imperfections of RMs, especially in complex and challenging tasks.
@@
Table 1: Accuracy improvement for each $\lambda$ on all the benchmarks tested at 200 step in PPO. Baseline here is the SFT model, which is LLaMa3-8B-Instruct (Dubey et al., 2024).
@@
Table 2: Accuracy improvement for each $\lambda$ on all the benchmarks tested at 800 step in PPO. Baseline here is the SFT model, which is LLaMa3-8B-Instruct (Dubey et al., 2024).
@@
Overall, the comparison shows that while the over-scoring strategy can occasionally provide short-term performance boosts, particularly in context-heavy tasks, the under-scoring (minimum reward) setting generally yields more reliable and stable improvements across a wider range of benchmarks. This underscores the value of conservative reward modeling, especially when dealing with complex reasoning tasks or when the RM itself may be noisy or imperfect. The results further validate the robustness of the reward-robust RLHF approach, where minimizing the reward signal helps mitigate the risk of performance degradation over time.
@@
C.3 Ablation study on BRME
@@
In this section, we present ablation study results on the RMs to highlight the superiority of BRME over the traditional RM. The coverage and margin advantages are detailed in Appendix B.3. Here, we provide a comparison of accuracy and the direct effect on the PPO process.
We first compared the accuracy of the single-head RM trained with MLE loss to the nominal head of the BRME on the preference dataset. The results in Table 5 clearly demonstrate that BRME achieves higher accuracy across most benchmarks, indicating a better ability to distinguish between preferred and non-preferred responses. Specifically, BRME outperforms the traditional RM by a notable margin across different categories such as General QA, Writing, Comprehension, and Math. The overall accuracy of BRME reaches 75.6%, compared to 74.5% for the traditional RM, showing that BRME’s ability to model uncertainty provides a tangible advantage in distinguishing between correct and incorrect responses.
@@
We also examined the impact of these RMs on the PPO process by comparing the performance of actors trained using BRME and the traditional RM. The results indicate that actors trained with BRME benefit from more stable and reliable reward signals, leading to smoother training curves and improved final performance. This is particularly evident in tasks requiring more nuanced decision-making, where BRME’s broader reward distribution prevents the actor from overfitting to narrow or incorrect reward signals.
@@ C.4 Ablation study on integration strategy
The mean strategy, while easier to implement, lacks the ability to differentiate between high-quality and low-quality reward signals effectively. By averaging all signals, it fails to account for the underlying uncertainty or variance in the RM. In contrast, the reward-robust RLHF approach, particularly with $\lambda=0.6$, preserves the benefits of more robust exploration and ensures that the model can continue improving in the later stages of training, as observed in the experiments.
@@
Figure 9: Ablation study on BRME and integration strategy.
@@
Appendix D Dataset Descriptions
Table 6 outlines the various datasets used for training BRME and the PPO models. These datasets span a wide range of task types, ensuring that the RM is exposed to diverse examples during training, helping it generalize across different domains. In UltraFeedBack, we selected response pairs where the score is greater than or equal to 2 to form the chosen and rejected response pairs, which were then used to train the BRME. By focusing on high-quality response pairs, we aim to ensure that the RM is robust and capable of accurately distinguishing between better and worse responses.
For evaluation, we utilized a comprehensive set of benchmarks, as detailed in Table 7. These benchmarks cover various domains, including general knowledge, logical reasoning, commonsense, and mathematical reasoning, allowing us to thoroughly assess the performance of the RMs across different types of tasks. To ensure consistency and reliability in our evaluation process, we employed the LM-evaluation-harness framework, which is a widely-used standard for evaluating LLMs (Gao et al., 2024). This framework provides a standardized and rigorous approach to comparing model performance across a diverse set of tasks, ensuring that the results are comparable with other works in the field.
By using a diverse set of training datasets and evaluation benchmarks, we aim to provide a comprehensive assessment of our model’s capabilities. The combination of varied training data ensures the robustness of the RM, while the evaluation benchmarks test the model’s ability to generalize across different domains, making our findings relevant to a wide range of applications.
Table 6: Description of the datasets in the training pipeline.
Datasets Size Description
Table 7: Description of the benchmarks.
Benchmark Size Description
TruthfulQA 817 A QA task aimed at evaluating the truthfulness and factual accuracy of model responses.
MMLU 28,128 Knowledge-based multi-subject multiple choice questions for academic evaluation.
GSM8K 1,000 A benchmark of grade school math problems aimed at evaluating reasoning capabilities.
FDA 551 Tasks for extracting key-value pairs from FDA documents to test information extraction.
Arithmetic 2,000 A small battery of 10 tests that involve asking language models a simple arithmetic problem in natural language.
ANLI 3,200 Adversarial natural language inference tasks designed to test model robustness. It collected via an iterative, adversarial human-and-model-in-the-loop procedure. It consists of three rounds that progressively increase in difficulty and complexity.
@@ Appendix E Proof for Lemma 1
Let the RM provide a constant reward $r(s,a)=c$ for all actions $a\in\mathcal{A}$ and states $s\in\mathcal{S}$ during PPO training. Then, the gradient of the PPO objective function with respect to the policy parameters $\theta$ is zero, implying that the actor cannot be optimized under such a RM.
@@
If the RM provides a constant reward $r(s_{t},a_{t})=c$ for all actions, the state-action value function and the state value function become:
@@ Appendix F Future works: Heterologous reward fusion
Heterologous Reward Fusion (HRF) aims to enhance the robustness and coverage of the uncertainty set in our RMs by incorporating multiple heterologous reward sources. This method involves combining several different RMs trained on diverse datasets, such as Baichuan2-33B (Yang et al., 2023), Qwen2-72B (Yang et al., 2024a), and LLaMa3-8B (Dubey et al., 2024). The key advantage of using heterologous models lies in the diversity of their training data, which produces more varied reward scores and helps to enrich the uncertainty set.
@@
However, one of the primary challenges in HRF is that each reward source has different value ranges, making direct comparisons potentially unfair. To address this, we perform empirical reward normalization. We score a separate dataset, HH-RLHF (Bai et al., 2022a), using each reward source and compute the mean and variance for each. During PPO training, we transform the rewards from each source by normalizing them based on their respective means and stds. This normalization helps ensure that the reward signals from different models are comparable, allowing for a more fair integration of the rewards.
Table 8 presents the results of our HRF experiment, showing the reward distribution characteristics (max, min, mean, and std) for each model, along with their accuracy. Accuracy here is defined as the proportion of instances where the model correctly identifies the chosen response as superior to the rejected one.
However, it is important to note that we did not conduct end-to-end PPO experiments incorporating the full RM pipeline in this exploration. The primary reason for this limitation is the significant computational resources and time required to carry out such experiments comprehensively. Running full-scale PPO experiments, especially when integrating multiple heterologous RMs, is computationally expensive and requires extended periods of training, particularly when dealing with large models like Baichuan2-33B and Qwen2-72B. Moving forward, our next steps include performing the full end-to-end PPO experiments to evaluate the impact of HRF on the performance of the trained policy.
Table 8: Reward distribution characteristics in several RMs trained from different base model.
@@
Appendix G Limitations
While the proposed reward-robust RLHF framework shows improved performance on automatic evaluation benchmarks, several non-negligible limitations remain. First, for some prompts, if all reward heads in the BRME fail to provide correct signals and exhibit similar error patterns, the reward hacking behavior seen in standard RLHF cannot be entirely mitigated, even with the reward-robust approach. This suggests that the ultimate optimization still heavily depends on the model’s capacity and generalization ability, particularly in handling out-of-distribution (OOD) data. As a result, improving the quality of the reward model and the data used to train PPO is likely to be the most critical factor influencing long-term training success.
Second, some of the theoretical hypotheses proposed to explain the experimental results in Section 6.1 still require further validation. Conducting fine-grained analysis, particularly regarding the specific impact of over-scoring and under-scoring signals on the PPO exploration process, will be highly valuable. Addressing these gaps will be one of our future research focuses.